# Automated Bug Generation in the era of Large Language Models

### v1

@@ Abstract.
Bugs are essential in software engineering; many research studies in the past decades have been proposed to detect, localize, and repair bugs in software systems. Effectiveness evaluation of such techniques requires complex bugs, i.e., those that are hard to detect through testing and hard to repair through debugging. From the classic software engineering point of view, a hard-to-repair bug differs from the correct code in multiple locations, making it hard to localize and repair. Hard-to-detect bugs, on the other hand, manifest themselves under specific test inputs and reachability conditions. These two objectives, i.e., generating hard-to-detect and hard-to-repair bugs, are mostly aligned; a bug generation technique can change multiple statements to be covered only under a specific set of inputs. However, these two objectives are conflicting for learning-based techniques: A bug should have a similar code representation to the correct code in the training data to challenge a bug prediction model to distinguish them. The hard-to-repair bug definition remains the same but with a caveat: the more a bug differs from the original code (at multiple locations), the more distant their representations are and easier to be detected.
@@
We propose BugFarm, to transform arbitrary code into multiple complex bugs. BugFarm leverages Large Language Models (LLMs) to mutate code in multiple locations (hard-to-repair). To ensure that multiple modifications do not notably change the code representation, BugFarm analyzes the attention of the underlying model and instructs LLMs to only change the least attended locations (hard-to-detect). Our comprehensive evaluation of $320k+$ bugs from over $2.5M$ mutants generated by BugFarm and two alternative approaches demonstrates our superiority in generating bugs that are hard to detect by learning-based bug prediction approaches (up to 41% higher False Negative Rate and 11%, 6%, 29%, and 21% lower Accuracy, Precision, Recall, and F1 score) and hard to repair by state-of-the-art learning-based program repair technique (22% repair success rate compared to 34% and 49% of LEAM and $\mu$BERT bugs). BugFarm is efficient, i.e., it takes nine seconds for it to mutate a code without any prior training time.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXcopyright: noneprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Problem Statement and Significance. Machine learning is becoming the new automation technology, and software engineering automation is no exception. AI pair programming techniques such as Copilot (GitHub, 2023) are integrated into IDEs, and LLMs such as GPT-4 (OpenAI, 2023) and Bard (Google, 2023a) are now accessible through API. Following this trend, state-of-the-art software analysis techniques either fine-tune pre-trained code-language models or prompt LLMs to automate code-related tasks  [^1] We differentiate between LLMs and code-language models. The former models are large ($>10B$ parameters) and are used through zero- or few-shot prompting. The latter models are smaller and are used after fine-tuning.. This demands automated techniques for generating high-quality datasets to assess their true effectiveness. Concerning bug-related tasks such as bug prediction, localization, and repair, such datasets should include a diverse set of complex bugs, i.e., those that are hard to detect to challenge bug prediction techniques and hard to repair for debugging approaches.
From the classic software engineering perspective, hard-to-detect bugs often exist at locations reachable only under a particular combination of test inputs or edge cases (Lu et al., 2005; Böhme and Roychoudhury, 2014). A similar definition does not stand concerning learning-based bug prediction techniques. That is because they do not care what inputs trigger the bug in the given code. In contrast, they look for bug patterns in their training/fine-tuning data (Yang et al., 2015) or to check if the code representation is closer to buggy or correct examples they have seen (Ibrahimzada et al., 2022). As a result, to challenge the learning-based bug prediction techniques, one should inject unseen bug patterns such that the code representation of the generated bug and correct code are similar.
Hard-to-repair bugs usually involve multiple statements, making debugging techniques struggle to localize and understand their interactions (Li et al., 2018) to fix them automatically. A similar definition holds from the learning-based software engineering, but with a slight consideration: The power of machine learning is in learning from patterns that persist in their dataset. Consequently, modifying a code in multiple locations, where the learning-based debugging techniques specifically look to localize or repair the bug, would be counter-intuitive. Can existing techniques provide complex bugs, as defined, to evaluate learning-based techniques?
Research Gap. Several attempts have been made to construct real-world bug datasets, e.g., Defects4J (Just et al., 2014a), BugSwarm (Tomassi et al., 2019), BugsInPy (BugsInPy, 2023), RegMiner (Song et al., 2022), and ManySStuBs4J (sst, 2023). While representing real-world bugs, such datasets are only limited to the projects and bug patterns included in their mining efforts. More importantly, they may be included in the training/fine-tuning data of learning-based techniques, causing data leakage (Samala et al., 2020). To overcome these limitations, researchers have proposed automated bug generation techniques (Tian et al., 2022; Khanfir et al., 2023; Jia and Harman, 2009). Such techniques, by default, change a few locations in the code (easy to repair and detect) but can be configured to modify multiple lines. That said, changing multiple lines randomly (Khanfir et al., 2023) or following a set of heuristics (Tian et al., 2022) may result in mutants with a different code representation than the correct code, making them easy to detect. Furthermore, changing lines that learning-based repair techniques have been trained to look into and repair (Xia et al., 2023), such as conditional statements, results in bugs that can be easily repaired, even if they involve multiple lines.
Proposed Solution. To advance automated bug generation concerning the evaluation of learning-based bug-related tasks, we propose BugFarm. For a given arbitrary code, BugFarm prompts a Large Language Model (LLM) to mutate multiple statements. Given that LLMs are potentially creative in generative tasks, leveraging them helps avoid overfitting to a limited number of mutation operators and bug patterns. To ensure that changing too many locations does not drastically change the code representation of the mutant, BugFarm analyzes the attention of the underlying model and instructs LLMs to only change those the model attends least to. As a result, the generated code has a similar representation to the original one but differs within multiple locations. This makes them hard to detect by bug prediction approaches and hard to repair by debugging techniques. BugFarm is programming language-agnostic, i.e., making it a pragmatic approach to generate bugs for any programming language. Like related mutant generation techniques, BugFarm creates unconfirmed bugs, which should be validated through test execution. Our notable contributions are:
  • A new perspective into bug generation: We propose a novel technique for bug generation concerning the evaluation of the learning-based bug-related tasks. Our proposed technique, BugFarm, relies on the model’s attention to different parts of code to identify where to inject bugs so that the impact of changes on code representation is minimal. It leverages such information to craft contextual prompts for LLMs to mutate any arbitrary code. The implementation and artifact of BugFarm, including all the generated bugs, are publicly available (authors, 2023).
  • Empirical evaluation: We extensively evaluated BugFarm and two most recent learning-based bug generation approaches, LEAM (Tian et al., 2022) and $\mu$BERT (Khanfir et al., 2023). Our evaluation of $320k+$ bugs (from over $2.5M$ mutants) generated for $15$ Java projects answers research questions related to (1) the properties of generated bugs, (2) the ability of learning-based bug prediction models to detect them, (3) the effectiveness of a learning-based program repair technique to repair them, (4) the impact of prompt crafting, and (5) the performance characteristics of techniques. The results confirm that BugFarm bugs are hard-to-detect (up to 41% higher False Negative Rate and 11%, 6%, 29%, and 21% lower Accuracy, Precision, Recall, and F1 score) and hard-to-repair (22% repair success rate compared to 34% and 49% of LEAM and $\mu$BERT bugs). Also, generating effective bugs is more efficient using BugFarm.
2. Illustrative Example
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(a) Original Method
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if ((bag == null)) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(b) Bug generated by LEAM (Tian et al., 2022)
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        new TransformedSortedBag<>(bag, transformer);
    if (!!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return dec;
(c) Bug generated by $\mu$BERT (Khanfir et al., 2023)
public static <E> TransformedSortedBag<E> transformedSortedBag(
        final SortedBag<E> bag,
        final Transformer<? super E, ? extends E> transformer) {
    final TransformedSortedBag<E> dec =
        null;
    if (!bag.isEmpty()) {
        @SuppressWarnings(”unchecked”)
        final E[] values = (E[]) bag.toArray();
        bag.clear();
        for (final E value : values) {
            dec.decorated().add(transformer.transform(value));
    return null;
(d) Bug generated by BugFarm
@@ 2. Illustrative Example
Figure 1. An illustrative example showing the bugs generated for the original code (a) using LEAM (b), $\mu$BERT (c), and BugFarm (d).
@@
To illustrate the limitations of prior work and present the key idea behind BugFarm, we will use the code snippets in Figure 1. The original code in Figure 0(a) takes bag and transformer objects as inputs. If bag is not empty, it moves all the objects inside it to transformer. There are several ways to inject bugs into this code snippet by adding, deleting, or modifying the $13$ statements in its body (some statements are split into two lines for better presentation). These bugs are intended to challenge bug prediction and program repair models we used in our evaluation (section 5), which are fine-tuned on top of CodeT5 (Wang et al., 2021). Concerning the attention of the underlying model to the original code, the two least attended statements (details in algorithm 1) are highlighted in green.
@@
We leverage three bug generation tools, namely, BugFarm, LEAM (Tian et al., 2022), and $\mu$BERT (Khanfir et al., 2023) to transform the original code snippets into bugs (bugs were confirmed through test execution). LEAM is a technique that represents code as a sequence of AST nodes and learns to apply grammar rules to select and modify the code for bug generation. $\mu$BERT selects code tokens corresponding to AST nodes, replaces them with a special token <mask>, and asks CodeBERT to replace them with new tokens for bug generation. BugFarm identifies the least attended statements and prompts an LLM to perform bug-inducing transformations only on those lines. Figures 0(b) and  0(c) show one of the generated bugs by LEAM and $\mu$BERT (LEAM and $\mu$BERT generate $19$ and $68$ mutants in total, out of which only $6$ and $1$ are confirmed bugs). BugFarm generates a bug shown in Figure 0(d). The lines changed to introduce a bug are highlighted in red.
@@
BugFarm bug is notable from various perspectives. First, it involves multiple statements (hard to localize and repair). Second, both modified statements are among the least attended statements, making it hard for the model to distinguish the bug from the original code (hard to detect). LEAM and $\mu$BERT bugs modify only one line that is not among the least attended statements. As a result, when running against our studied bug prediction model (section 5.3) and program repair model (section 5.4), they were easily detected and repaired. In contrast, the same models failed to detect and repair BugFarm bug.
Figure 2. Overview of BugFarm
3. Approach Overview
@@
Figure 2 provides an overview of BugFarm framework consisting of three major components: (1) Method Extractor, (2) Attention Analyzer, and (3) Bug Generator. BugFarm takes a project and a pre-trained code-language model as inputs and extracts the methods through a lightweight static analysis. For each method, it then identifies the bug location candidates by analyzing the model’s attention to individual code tokens. To generate bugs for each method, BugFarm crafts a prompt to an LLM, including the code to be transformed into bugs and additional contexts reflecting candidate locations for bug injection. From the generated bugs, BugFarm selects those that are syntactically correct, where the bug-inducing changes do not significantly change the attention of the model.
@@
The Method Extractor component takes the input project and builds its corresponding parse tree to extract all the methods in the source files. These methods will be passed as an input to Attention Analyzer to identify the candidate locations for bug injection (more details in  section 4.1).
@@
Recent state-of-the-art learning-based software engineering models are all transformer-based due to the naturalness of code (Hindle et al., 2016) and the promising results of such models in natural language processing (NLP) (Vaswani et al., 2017). Transformer-based models rely on attention for neural code representation. So, to identify the candidate locations to inject unnoticeable bugs into a given method, Attention Analyzer extracts the corresponding attention of the model to code tokens and identifies those with the lower contribution in the code representation, i.e., those with less attention weight values. Bug locations in BugFarm are at the statement level; thereby, BugFarm ranks the input method’s statements based on the #$\#T_{LA}/\#T$ values and chooses the ones with the lowest value. Here, $T_{LA}$ is the number of least attended tokens, and $T$ denotes the total number of tokens in the statement (more details in  section 4.2).
@@
Finally, Bug Generator component takes the list of bug location candidates as input and creates a prompt consisting of natural language instruction, bug location candidates, and the method. It then sends the prompt to an LLM and collects the generated bugs in the response. For the generated bugs by LLM, BugFarm computes the extent to which the bug-inducing changes impact the model’s attention to the buggy method compared to the input method, and selects those with negligible impact on the attention. It also discards the bugs that are syntactically incorrect for the sake of quality (more details in section 4.3).
4. BugFarm
@@
In this section, we will explain the details of three main components of BugFarm. Specifically, we will first discuss how BugFarm parses the subjects and extract methods as inputs for the Attention Analyzer component. Next, we will provide a background on the attention analysis required to understand the subsequent components. Finally, we answer two main questions, namely (1) How does BugFarm decide where to inject bugs?, and (2) how BugFarm generates a set of bugs?
4.1. Method Extractor
Our approach sets itself apart from most prior work, which relies on defect datasets for automated bug generation (Tian et al., 2022; Nong et al., 2023; Patra and Pradel, 2021). Instead, it focuses on injecting diverse and complex bugs into any well-maintained project using LLMs, regardless of the programming language. As a result, we assume that the input to the BugFarm is a project rather than a single method. As a first step, BugFarm first requires extracting the implemented methods in each input project. To that end, Method Extractor component leverages a language-specific parser depending on the programming language used in the input project, builds program parse trees, and extracts a list of methods and constructors. Next, it collects both signatures and body (excluding docstrings) using MethodDeclaration and ConstructorDeclaration blocks in the parse tree. Each method will be passed to the Attention Analyzer component, which we will describe next.
4.2. Attention Analyzer
@@
State-of-the-art code-language models are based on the Transformer architecture (Vaswani et al., 2017)—an encoder to capture input’s contextual representations and a decoder to generate output tokens. To produce contextualized vector representation of a sequence of tokens, Transformers rely on Multi-Head Self-Attention. More specifically, for a method that consists of $n$ tokens $\overrightarrow{Tkn}=\{m_{0},\ldots,m_{n-1}\}$, a Transformer model with $\mathcal{L}$ layers takes $\overrightarrow{Tkn}$ as input and produces $\textbf{H}^{\ell}=[\textbf{h}_{0}^{\ell},...,\textbf{h}_{n-1}^{\ell}]$. Here, $\textbf{H}^{\ell}$ corresponds to the hidden vector representations in layer $\ell\in\{1,2,...,\mathcal{L}\}$. In each Transformer layer $\ell$, multiple self-attention heads are used to aggregate the outputs of the previous layer. Consequently, for each token $m_{i}$ in the method, the self-attention assigns a set of attention weights with respect to all other tokens in the input sequence, i.e., $Attention(m_{i})=\{\alpha_{i0},\dots,\alpha_{in-1}\}$, where $\alpha_{ij}$ indicates the relative attention of $m_{i}$ to $m_{j}$.
@@
The hidden representation weights of Transformers will be computed based on the attention weights, making the attention weights crucial to identify the importance of each token in the final representation of the method. Recent research has shown that attention analysis can be utilized to interpret the knowledge learned during training (Hernández López et al., 2022; Wan et al., 2022; Ibrahimzada et al., 2022). Following their insights, BugFarm leverages self-attention weight analysis to scrutinize the input method and identify tokens (and subsequently, statements) with the lowest attention weights. These statements are where BugFarm can change during bug-inducing transformation without impacting the overall representation of the method. Algorithm 1 explains our approach for attention analysis, which takes the method to be transformed into a bug, threshold value $k$, and a transformer-based model as inputs and pinpoints the $k\%$ of the least attended statements $LAS$ as outputs. To that end, it first extracts the list of least attended tokens in the method $LAT$ (Lines 1-10) and uses them to pinpoint the least attended statements $LAS$ (Lines 11-20).
@@
With the least attended tokens extracted, the algorithm can identify the least attended statements, $LAS$. To that end, it weighs each statement by a score (Lines 11-14), which is the ratio of the number of least attended tokens in that statement, normalized by its length. The key idea here is that a statement with the highest amount of overlap between least attended tokens should achieve a lower score, thus, be considered the least attended statement. Without normalizing the statement length, longer statements will be penalized, and BugFarm never selects them as a bug injection location. Finally, the statements are sorted in ascending order based on their scores, and the least $k\%$ attended statements (we take the ceiling in case $k\%$ of total statements is less than one) will be returned as $LAS$ (Lines 15-20).

### v2

@@ Abstract
Bugs are essential in software engineering; many research studies in the past decades have been proposed to detect, localize, and repair bugs in software systems. Effectiveness evaluation of such techniques requires complex bugs, i.e., those that are hard to detect through testing and hard to repair through debugging. From the classic software engineering point of view, a hard-to-repair bug differs from the correct code in multiple locations, making it hard to localize and repair. Hard-to-detect bugs, on the other hand, manifest themselves under specific test inputs and reachability conditions. These two objectives, i.e., generating hard-to-detect and hard-to-repair bugs, are mostly aligned; a bug generation technique can change multiple statements to be covered only under a specific set of inputs. However, these two objectives conflict in the learning-based techniques: A bug should have a similar code representation to the correct code in the training data to challenge a bug prediction model to distinguish them. The hard-to-repair bug definition remains the same but with a caveat: the more a bug differs from the original code (at multiple locations), the more distant their representations are and easier to detect. This demands new techniques to generate bugs to complement existing bug datasets to challenge learning-based bug prediction and repair techniques.
@@
We propose BugFarm to transform arbitrary code into multiple hard-to-detect and hard-to-repair bugs. BugFarm mutates code in multiple locations (hard-to-repair) but leverages attention analysis to only change the least attended locations by the underlying model (hard-to-detect). Our comprehensive evaluation of $435k+$ bugs from over $1.9M$ mutants generated by BugFarm and two alternative approaches demonstrates our superiority in generating bugs that are hard to detect by learning-based bug prediction approaches (up to 40.53% higher False Negative Rate and 10.76%, 5.2%, 28.93%, and 20.53% lower Accuracy, Precision, Recall, and F1 score) and hard to repair by state-of-the-art learning-based program repair technique (28% repair success rate compared to 36% and 49% of LEAM and $\mu$BERT bugs). BugFarm is efficient, i.e., it takes nine seconds to mutate a code with no training overhead.
I Introduction
@@
Machine learning is the new automation technology, and software engineering is no exception: State-of-the-art software analysis techniques either fine-tune pre-trained models or prompt Large Language Models (LLMs) to automate code-related tasks. This demands the existence of high-quality datasets to assess their actual effectiveness. Concerning bug-related tasks, such datasets should include a diverse set of complex bugs, i.e., those that are hard to detect to challenge bug prediction techniques and hard to repair for debugging approaches.
@@
Figure 1: An example showing the bugs generated for the original code (a) using LEAM (b), $\mu$BERT (c), and BugFarm (d)
From the classic software engineering perspective, hard-to-detect bugs often exist at locations reachable only under a particular combination of test inputs or edge cases [27, 7]. However, a similar definition does not stand concerning learning-based bug prediction techniques. That is because they do not care what inputs trigger the bug in the given code. In contrast, they look for bug patterns in their training/fine-tuning data [46] or to check if the code representation is closer to buggy or correct examples they have seen [20]. As a result, existing real-world bug datasets such as Defects4J [23], BugSwarm [40], BugsInPy [10], RegMiner [37], and ManySStuBs4J [2], while necessary to assess their effectiveness in dealing with real-world bugs, are not enough to assess their true effectiveness and rule out data contamination [35]: To challenge the learning-based bug prediction techniques, one should inject unseen bug patterns such that the code representation of the generated bug and correct code are similar.
Hard-to-repair bugs usually involve multiple statements to challenge debugging techniques, localizing [26] and fixing automatically. A similar definition is held in learning-based software engineering, but with slight consideration: Modifying a code in multiple locations can change the representation of buggy code, making it easy for bug prediction models to detect it. These two conflicting objectives make the problem of complex bug generation for learning-based techniques challenging. Existing automated mutant generation techniques [39, 24, 21], by default, change a few locations in the code, making their bugs easy to repair and detect. Some can be configured to modify multiple lines randomly [24] or follow a set of heuristics [39]. However, this may result in mutants with a different code representation than the correct code, making them easy to detect.
To advance automated bug generation concerning the evaluation of learning-based bug-related tasks, we propose BugFarm. For a given code, BugFarm prompts an LLM to mutate multiple statements. Given that LLMs are potentially creative in generative tasks, leveraging them helps avoid overfitting to a limited number of mutation operators and bug patterns. To ensure that changing many locations does not drastically impact the code representation of mutants, BugFarm analyzes the attention of the underlying model and instructs LLMs only to change those the model attends least to. The generated code has a similar representation to the original one. Still, it differs within multiple locations, making them hard to detect by bug prediction approaches and hard to repair by debugging techniques. BugFarm is language-agnostic and can generate bugs for any programming language. Like alternative approaches, the BugFarm creates unconfirmed bugs and requires a validation process to confirm them as bugs. Our notable contributions are:
  • A new perspective into synthetic bug generation: We propose a novel technique for bug generation concerning the evaluation of the learning-based bug-related tasks. We do not claim our bugs challenge classic bug detection (testing) techniques, as we propose a new definition for hard-to-detect bugs. Our goal is to demonstrate how ignoring properties of learning-based techniques in mutant/bug generation results in subpar assessment of them. The implementation of BugFarm and all the generated bugs are publicly available [4].
  • Empirical evaluation: We extensively evaluated BugFarm and two most recent learning-based mutant generation approaches, LEAM [39] and $\mu$BERT [24]. Our evaluation of $435k+$ bugs (from over $1.9M$ mutants) generated for $15$ Java projects confirm that compared to alternative approaches, BugFarm bugs are hard-to-detect (up to 40.53% higher False Negative Rate and 10.76%, 5.2%, 28.93%, and 20.53% lower Accuracy, Precision, Recall, and F1 score) and hard-to-repair (28% repair success rate compared to 36% and 49% of LEAM and $\mu$BERT bugs) by learning-based techniques. A large-scale human study with $97$ participants and $500$ sampled survived mutants from these techniques shows that BugFarm mutants are harder to decide buggy or equivalent. Also, generating effective bugs is more efficient using BugFarm.
II Illustrative Example
@@
To illustrate the limitations of prior work and present the key ideas behind BugFarm, we use the code snippets in Figure 1. There are several ways to inject bugs into original code by adding, deleting, or modifying the $11$ statements in its body (some statements are split into two lines for better presentation). However, the highlighted lines in green show the two least attended statements (details in algorithm 1).
Figure 2: Overview of BugFarm
@@
Figures 1b–1d show the bugs generated by three mutant generation tools, BugFarm, LEAM [39], and $\mu$BERT [24] (bugs were confirmed through test execution). LEAM considers code as a sequence of AST nodes and learns to apply grammar rules to select and modify the code for bug generation. $\mu$BERT selects code tokens corresponding to AST nodes, replaces them with a special token <mask>, and asks CodeBERT to replace them with new tokens for bug generation. BugFarm identifies the least attended statements and prompts an LLM to perform bug-inducing transformations only on those lines. LEAM and $\mu$BERT generate $19$ and $68$ mutants in total, out of which only $6$ and $1$ are confirmed bugs, respectively. BugFarm generates the bug in Figure 1d. The lines changed to introduce a bug are highlighted in red.
@@
BugFarm bug is notable from various perspectives. First, it involves multiple statements (hard to localize and repair). Second, both modified statements are among the least attended statements, making it hard for the model to distinguish the bug from the original code (hard to detect). LEAM and $\mu$BERT bugs modify only one line that is not among the least attended statements. As a result, when running against our studied bug prediction model (section V-C) and program repair model (section V-D), they were easily detected and repaired. In contrast, the same models failed to detect and repair BugFarm bug.
III Approach Overview
@@
Figure 2 provides an overview of BugFarm framework consisting of three major components: (1) Method Extractor, (2) Attention Analyzer, and (3) Bug Generator. BugFarm takes a project and a pre-trained code-language model as inputs and extracts the methods through a lightweight static analysis. For each method, it identifies the change location candidates by analyzing the model’s attention to individual code tokens. To generate mutants for each method, BugFarm crafts a prompt to an LLM, including the original code and additional contexts reflecting candidate locations for mutations.
@@
The Method Extractor component takes the input project and builds its corresponding parse tree to extract all the methods in the source files. These methods will be passed as an input to Attention Analyzer to identify the candidate locations for the mutation (section IV-A).
@@
Recent state-of-the-art learning-based software engineering models are all transformer-based, which rely on attention for neural code representation. So, to identify the candidate locations to inject unnoticeable bugs, Attention Analyzer extracts the corresponding attention of the model to code tokens and identifies those with the lower contribution in the code representation, i.e., those with less attention weight values. Bug locations in BugFarm are at the statement level; thereby, BugFarm ranks the input method’s statements based on the #$\#T_{LA}/\#T$ values and chooses the ones with the lowest value. Here, $T_{LA}$ is the number of least attended tokens, and $T$ denotes the total number of tokens in the statement (section IV-B).
@@
Finally, Bug Generator component takes the list of location candidates and creates a prompt consisting of natural language instruction, location candidates, and the original code. For the generated mutants, BugFarm computes the extent to which the changes impact the model’s attention, and selects mutants with negligible impact on the code representation. It also discards the syntactically incorrect bugs for the sake of quality (section IV-C).
IV BugFarm
@@
In this section, we first discuss how BugFarm parses the subjects and extracts methods as inputs for the Attention Analyzer component. Next, we will provide a background on the attention analysis required to understand the subsequent components. Finally, we answer two main questions, namely (1) How does BugFarm decide where to inject bugs?, and (2) how BugFarm generates a set of bugs?
IV-A Method Extractor
BugFarm can take a single method or a repository as input. In the latter case, it first needs to extract the implemented methods in the project for further mutation. To that end, Method Extractor component leverages a parser depending on the programming language used in the input project, builds program parse trees, and extracts a list of methods and constructors. Next, it collects the signature and body of the method (excluding docstrings) and passes them to the Attention Analyzer component.
IV-B Attention Analyzer
@@
State-of-the-art code-language models are based on the Transformer architecture [42]. To produce contextualized vector representation of a sequence of tokens, Transformers rely on Multi-Head Self-Attention. For a method that consists of $n$ tokens $\overrightarrow{Tkn}=\{m_{0},\ldots,m_{n-1}\}$, a Transformer model with $\mathcal{L}$ layers takes $\overrightarrow{Tkn}$ as input and produces $\textbf{H}^{\ell}=[\textbf{h}_{0}^{\ell},...,\textbf{h}_{n-1}^{\ell}]$. Here, $\textbf{H}^{\ell}$ corresponds to the hidden vector representations in layer $\ell\in\{1,2,...,\mathcal{L}\}$. In each Transformer layer $\ell$, multiple self-attention heads are used to aggregate the previous layer’s outputs. Consequently, for each token $m_{i}$, the self-attention assigns a set of attention weights concerning all other tokens in the input sequence, i.e., $Attention(m_{i})=\{\alpha_{i0},\dots,\alpha_{in-1}\}$, where $\alpha_{ij}$ indicates the relative attention of $m_{i}$ to $m_{j}$.
@@
Attention weights reflect the importance of each token in the final code representation. Hence, BugFarm analyze attention weights to identify tokens (and subsequently, statements) with the lowest attention weights. These statements are where BugFarm can change during bug-inducing transformation without impacting the overall representation of the method. Algorithm 1 explains our approach for attention analysis, which takes the original method, a threshold value $k$, and a transformer-based model as inputs and pinpoints the $k\%$ of the least attended statements $LAS$ as outputs. To that end, it first extracts the list of least attended tokens in the method $LAT$ (Lines 1-10) and uses them to pinpoint the least attended statements $LAS$ (Lines 11-20).
@@
1
2